OpenStudy (anonymous):

If h(x)=(f o g) (x) and h(x)=3√x+3, find g(x) if f(x)=3√x+2? The 3 is in top the the square root. Please help me! :(

OpenStudy (zzr0ck3r):

if h(x) = g(x) then putting in f(x) = x will make g(x) = h(x) = g(f(x))

OpenStudy (anonymous):

@zzr0ck3r so then would it be 3sqrt x+3+2.....?

OpenStudy (zzr0ck3r):

are h(x) and g(x) the same thing?

OpenStudy (anonymous):

@zzr0ck3r no

OpenStudy (larseighner):

\( h(x) = (f\circ g)(x) = f\left(g(x)\right) \)

OpenStudy (larseighner):

Now \( h(x) = 3\sqrt x + 3 \) and \(f(x) = 3\sqrtx + 2 \) so the question is what g(x) will turn f(x) into h(x)?

OpenStudy (anonymous):

@LarsEighner so would the overall answer be? I've done it more than once and keep getting different answers

OpenStudy (larseighner):

workin on it

OpenStudy (anonymous):

@LarsEighner ok

OpenStudy (larseighner):

okay I think this is it we want u such that \( 3\sqrt x + 3 = 3 \sqrt u + 2 \)

OpenStudy (larseighner):

which is: \(\begin{align} 3\sqrt x +3 - 2 &= 3\sqrt u \cr \sqrt x - {1\over 3} &= \sqrt u \cr (\sqrt x + 1)^2 = u \end{align}\)

OpenStudy (anonymous):

@LarsEighner I thought the outcome was just a plain number, guess I was wrong. Thank you.

OpenStudy (larseighner):

so \( g(x) = (\sqrt x + {1 \over 3})^2 \) check \( f(g(x)) = 3 \sqrt{(\sqrt x + {1\over 3})^2} + 2 \) \( f(g(x)) = 3 \sqrt x + (3){1\over 3} + 2 \) ( f(g(x)) = 3 \sqrt x + 3 \) which, lo an behold is the same as \( h(x) = 3 \sqrt x + 3 \)

OpenStudy (anonymous):

@LarsEighner Thank you very much. I kept making simple mistakes, I see what I did wrong now.